OpenStudy (anonymous):

A theater has a seating capacity of 663 and charges $4 for children, $6 for students, and $8 for adults. At a certain screening with full attendance, there were half as many adults as children and students combined. The receipts totaled $4016. How many children attended the show?

OpenStudy (alexwee123):

x=children, y=student, z=adult 4x+6y+8z=4016 x+y+z=663 2z=x+y =>x+y-2z

OpenStudy (alexwee123):